Which of the expressions is not equal to the other three?
A. 7[a +2]
B. 20 + 8a -6 -a
C.17 +3[a -1] + 4a
D. 2a + 5[a -1] + 8

Found 3 solutions by greenestamps, ikleyn, Edwin McCravy:
Answer by greenestamps(13200)



Simplify each expression by combining like terms and/or using the distributive property to get rid of parentheses.

A. 7(a+2) = 7(a)+7(2) = 7a+14
B. 20+8a-6-a = (8a-a)+(20-6) = 7a+14
C. 17+3(a-1)+4a = 17+3(a)+3(-1)+4a = 17+3a-3+4a = (3a+4a)+(17-3)=7a+14
D. 2a+5(a-1)+8 = 2a+5(a)+5(-1)+8 = 2a+5a-5+8 = 7a+3

ANSWER: D


Answer by ikleyn(52788)

.
Which of the expressions is not equal to the other three?
A. 7[a +2]
B. 20 + 8a -6 -a
C.17 +3[a -1] + 4a
D. 2a + 5[a -1] + 8
~~~~~~~~~~~~~~~~~~~~~~~~~~~

One way to solve this problem is to simplify each expression to its simplest form
and compare them.


I will use another way. I will quickly calculate in my head the value of each expression for a = 0.


(A)  gives the value of 14.

(B)  gives the value of 14.

(C)  gives the value of 14.

(D)  gives the value of  3.


So, I conclude that expression (D) is not equal to (is different from) other three.

Solved.



Answer by Edwin McCravy(20056)

That sort of thing saves lots of time on such questions on exams.  If
substituting 0 gives the same results on some answers, substitute 1 or -1.

Edwin